# Feiyu Xiong3, Tong Xu2, Matthew B. Blaschko1

### v2

@@ Abstract
Large language models (LLMs) excel in generating coherent text, but they often struggle with context awareness, leading to inaccuracies in tasks requiring faithful adherence to provided information. We introduce FastMem, a novel method designed to enhance instruction fine-tuned LLMs’ context awareness through fast memorization of the prompt. FastMem maximizes the likelihood of the prompt before inference by fine-tuning only the last Feed-Forward Network (FFN) module. This targeted approach ensures efficient optimization without overfitting, significantly improving the model’s ability to comprehend and accurately follow the context. Our experiments demonstrate substantial gains in reading comprehension, text summarization and adherence to output structures. For instance, FastMem improves the accuracy of Llama 3-8B-Inst on the NQ-SWAP dataset from 59.1% to 71.6%, and reduces the output structure failure rate of Qwen 1.5-4B-Chat from 34.9% to 25.5%. Extensive experimental results highlight FastMem’s potential to offer a robust solution to enhance the reliability and accuracy of LLMs in various applications. Our code is available at: https://github.com/IAAR-Shanghai/FastMem.
@@
Junyi Zhuthanks: Equal Contribution1, Shuochen Liu2, Yu Yu3, Bo Tangthanks: Correspondence: tangb@iaar.ac.cn2, Yibo Yan4, Zhiyu Li3, Feiyu Xiong3, Tong Xu2, Matthew B. Blaschko1 1KU Leuven, 2 University of Science and Technology of China, 3 Institute for Advanced Algorithms Research, Shanghai, 4 National University of Singapore
@@ 1 Introduction
Despite their remarkable utility, LLMs have difficulty in fully understanding and following the provided context, which we refer as the ability of context awareness. A major aspect is maintaining faithfulness to the contextual information. For instance, in summarization tasks, LLMs may fabricate entities or relationships that are not present in the original context (Pagnoni et al., 2021; Cao et al., 2020; Maynez et al., 2020). This issue also reduces the effectiveness of Retrieval-Augmented Generation (RAG) (Liu et al., 2023; Chen et al., 2024), as the LLMs may falsify the information in references found by RAG when answering the question. Another aspect is adhering to a predefined output structure, which is necessary for many applications  (Liu et al., 2024).
Figure 1: Convention paradigm of LLM-based Question Answering Task with retrieved reference information (a), and our proposed FastMem method to enhance context awareness of LLMs (b).
@@
As illustrated in Figure 1(a), these findings imply that LLMs perform worse when they have high perplexity regarding the context. We also verify that models perform worse in extractive question answering when the perplexity of the answer span in the context is high (see Sec. 3.2). Inspired by these insights, we propose a novel approach named FastMem to improve context awareness via fast memorization of prompts, as depicted in Figure 1(b). Specifically, we adapt an instruction fine-tuned model by maximizing the prompt likelihood before generation. To ensure the efficiency of the optimization and prevent overfitting to small text snippets, we fine-tune only the last FFN module, achieving significant improvements with minimal computational cost. Our contributions are summarized as follows:
@@ 3 Method
In this section, we elaborate on our method, FastMem. In Sec. 3.1, we present the notations. In Sec. 3.2, we clarify the motivation behind our method. In Sec. 3.3, we explain how we develop our method to be robust and efficient. In Sec. 3.4, we explore combining FastMem with existing decoding strategies. Finally, in Sec. 3.5, we demonstrate the use cases. An illustration of our approach and its variants is provided in Fig. 2.
@@ 3.2 Motivation
As introduced in Sec. 1, conflicts between parametric knowledge and contextual information, as well as mass-seeking behavior are supposed to hamper the context awareness. To verify that, we compare the performance of Llama 3-8B-Instruct (Meta AI, 2023) on two datasets: Natural Questions (NQ) (Kwiatkowski et al., 2019) and NQ-SWAP. NQ is constructed based on Wikipedia, which is included in many pre-training corpora. Therefore, the information from NQ has likely been parameterized into Llama 3 (Touvron et al., 2023). NQ-SWAP is constructed from NQ using the method of Longpre et al. (2021), where we swap the answer strings, which are entities, with random similar entities from the corpus. This process creates new context-answer pairs that the model has not observed before.
We measure the model’s perplexity regarding the minimal span of tokens in the context containing the answer string. As shown in Fig. 3, we observe that the model exhibits higher perplexity on NQ-SWAP compared to NQ, proving that the information in NQ belongs to the model’s parametric knowledge. Additionally, the model suffers an obvious utility drop on NQ-SWAP, although the differences between the two datasets are merely short strings involving entities, indicating a negative correlation between perplexity and performance.
The negative correlation suggests that optimizing the model to memorize the context before inference could be beneficial. This approach can potentially address the commonly hypothesized causes of context-awareness issues. First, reducing the perplexity on context can alleviate potential conflicts between parametric knowledge and contextual information. Second, by increasing the model’s certainty regarding the text in the context, the mass-seeking effect will be reduced.
@@ 3.4 Integrating decoding strategies with FastMem
Li et al. (2023) propose contrastive decoding (CD) which considers the difference between the likelihoods predicted by a strong model and a weak model. Following the concept of CD, we make the prediction of FastMem contrasted with the original prediction made by the model before any memorization took place, as depicted in Figure 2(b). This is done by subtracting the original prediction probabilities $P_{\bm{\theta}_{0}}(\bm{x}_{t}|\bm{x}_{<t})$ from the current probabilities $P_{\bm{\theta}_{N}}(\bm{x}_{t}|\bm{x}_{<t})$. This subtraction aims to emphasize the differences introduced by the memorization process and mitigate any overfitting by highlighting where the model’s behavior has significantly changed. The modified prediction $\tilde{P}_{\bm{\theta}_{N}}(\bm{x}_{t}|\bm{x}_{<t})$ can be expressed as:
@@
In the second strategy, the final prediction is contrasted with an intermediate representation within the network, following the DoLa framework (Chuang et al., 2024). Specifically, we take the output from an earlier transformer block $k$ where $k<K$, and pass it through the final projection layer and softmax function to obtain an early-exit prediction $P_{\bm{\theta_{0}}^{\{1:k\}}}(\bm{x}_{t}|\bm{x}_{<t})$, as depicted in Figure 2(c). The final output $P_{\bm{\theta_{N}}}(\bm{x}_{t}|\bm{x}_{<t})$ is then adjusted by subtracting the early-exit prediction. This approach leverages the hierarchical nature of transformer models, where different layers capture different levels of abstraction. By contrasting the final output with an early-exit prediction, we can emphasize the knowledge learned and output by the last fine-tuned FFN. The adjusted prediction can be formalized as:
@@ 4.1 Experimental configurations
Our experiments include evaluations of popular open-source instruction fine-tuned LLMs from the Llama 3 series (8B and 70B) and the Qwen 1.5 series (4B, 7B, and 14B). Additionally, we compare our method with advanced decoding strategies: Decoding by Contrasting Layers (DoLa) (Chuang et al., 2024) and Context-Aware Decoding (CAD) (Shi et al., 2023).
@@
Prompts are carefully crafted and tailored to each dataset and task to elicit optimal performance of the original model, as detailed in Sec. B. For Q&A tasks, we prompt the model to perform extractive answering, while for summarization tasks, we prompt the model to output in required conciseness. Consistent templates are used across models to maintain uniformity in input formatting, with control tokens specialized for different models.
@@ 4.2 Improving context awareness
Tab. 1 presents the performance of various methods using Llama 3-8B-Instruct. Our method consistently outperforms the baseline and is competitive with or superior to decoding strategies across different datasets. In scenarios where there are conflicts between parametric knowledge and contextual information or where the instruction is counter-intuitive for the LLMs’ prior knowledge, our method significantly enhances the baseline model (NQ-SWAP: 59.1% $\rightarrow$ 71.6%, MemoTrap: 77.9% $\rightarrow$ 89.3%). For standard tasks, we find that FastMem is particularly beneficial for complex tasks such as multi-hop reading comprehension (HotpotQA: 67.1% $\rightarrow$ 69.3%). Furthermore, standalone FastMem is robust and generally improves the baseline model across tasks. In contrast, while CAD performs well on NQ-SWAP and MemoTrap, it is detrimental when applied to NQ and HotpotQA.
@@
Tab. 3 demonstrates the effectiveness and robustness of our method across various model architectures and sizes on Q&A tasks. For the Qwen series models, our method consistently enhances performance over the baseline on NQ-SWAP (7B-Chat: 71.9% $\rightarrow$ 75.2%, 14B-Chat: 69.9% $\rightarrow$ 75.6%) and MemoTrap (7B-Chat: 88.8% $\rightarrow$ 91.9%, 14B-Chat: 83.4% $\rightarrow$ 92.5%). We observe that FastMem marginally decreases performance on standard tasks (NQ and HotpotQA) for Qwen 1.5-7B-Chat, but this phenomenon is alleviated for Qwen 1.5-14B-Chat. In contrast, the improvements for Llama 3-70B-Instruct are significant and consistent across all datasets, which has also been observed for Llama 3-8B-Instruct, as shown in Tab. 1. This implies that larger and stronger models can benefit more from our method.
@@
In Q&A tasks, we design prompts to guide the model to output structured answers (see Sec. B), thereby preventing the model from blindly increasing accuracy by repeating given contexts. Structured output is also crucial for many applications of LLMs (Liu et al., 2024). However, we find that small-scale models like Qwen 1.5-4B-Chat may produce responses that do not adhere to the template. We mitigate this issue by incorporating format instructions into training, as shown in Fig. 5. Results are provided in Tab. 2. Our approach significantly reduces Qwen 1.5-4B-Chat’s failure rate in producing the required output format (NQ: 33.8% $\rightarrow$ 26.1%, NQ-SWAP: 34.9% $\rightarrow$ 25.5%), thereby greatly improving accuracy. Moreover, we observe that in-context learning (ICL) cannot easily address this issue, potentially because ICL examples extend the input length and make it more challenging for small-scale models to understand. Existing contrastive decoding strategies, such as CAD, also cannot address this problem.
@@ 4.3 Ablation study
FastMem is inspired by the negative correlation between perplexity and performance (see Sec. 3.2). To show that performance increases as perplexity decreases, we present the number of epochs vs. performance using Llama 3-8B-Instruct in Fig. 8. As we can see, performance improves as perplexity decreases up to 10 epochs. We also note that after 10 epochs, performance plateaus while perplexity continues to decrease, suggesting that the number of training epochs should not be too many.
@@ Appendix A Hyperparameters
Appendix B Prompt
Tab. 7 shows the prompts for Llama 3 across all datasets. For models of the Qwen 1.5 series, we simply replace the control tokens. Our prompts are carefully designed to guide the model to respond in alignment with the desired answers. For Llama 3-8B-Instruct, Llama 3-70B-Instruct, Qwen 1.5-8B-Chat, and Qwen 1.5-14B-Chat, we adopt the memorization prompt and optimize the model to memorize the reference, as shown in Fig. 4, in order to improve reading comprehension and summarization performance. To enhance the structured output of Qwen 1.5-4B-Chat, we adopt a memorization prompt that incorporates the instruction segment into memorization, as shown in Fig. 5.

### v3

@@ Abstract
Large language models (LLMs) excel in generating coherent text, but they often struggle with context awareness, leading to inaccuracies in tasks requiring faithful adherence to provided information. We introduce FastMem, a novel method designed to enhance instruction fine-tuned LLMs’ context awareness through fast memorization of the prompt. FastMem maximizes the likelihood of the prompt before inference by updating only the last Feed-Forward Network (FFN) module. This targeted approach ensures efficient optimization without overfitting, significantly improving the model’s ability to comprehend and accurately follow the context. Our experiments demonstrate substantial gains in reading comprehension, text summarization and adherence to output structures. For instance, FastMem improves the accuracy of Llama 3-8B-Inst on the NQ-SWAP dataset from 59.1% to 71.6%, and reduces the output structure failure rate of Qwen 1.5-4B-Chat from 34.9% to 25.5%. Extensive experimental results highlight FastMem’s potential to offer a robust solution to enhance the reliability and accuracy of LLMs in various applications. Our code is available at: https://github.com/IAAR-Shanghai/FastMem.
@@
Junyi Zhuthanks: Equal Contribution1, Shuochen Liu2, Yu Yu3, Bo Tangthanks: Correspondence: tangb@iaar.ac.cn2,3, Yibo Yan4, Zhiyu Li3, Feiyu Xiong3, Tong Xu2, Matthew B. Blaschko1 1ESAT-PSI, KU Leuven, 2 University of Science and Technology of China, 3 Institute for Advanced Algorithms Research, Shanghai, 4 National University of Singapore
@@ 1 Introduction
Despite their remarkable utility, LLMs have difficulty in fully understanding and following the provided context, which we refer as the ability of context awareness. In this paper, we consider both the instruction in the prompt, such as query or predefined output format, and the reference provided by the user or retrieval system as the context. A major issue with context awareness is maintaining faithfulness to the contextual information. For instance, in summarization tasks, LLMs may fabricate entities or relationships that are not present in the original reference (Pagnoni et al., 2021; Cao et al., 2020; Maynez et al., 2020). This issue also reduces the effectiveness of Retrieval-Augmented Generation (RAG) (Liu et al., 2023; Chen et al., 2024), as the LLMs may falsify the information in references found by RAG when answering the question. Another aspect is adhering to a predefined output structure, which is necessary for many applications  (Liu et al., 2024).
Figure 1: An illustrative example where an LLM cannot follow the reference due to its conflict with the LLM’s prior knowledge (a). In such cases, LLM often exhibits high perplexity on the conflicted information. Our proposed FastMem addresses this issue by enabling the model to memorize the reference text (thereby reducing perplexity) before conducting inference (b).
@@
These findings imply that LLMs perform worse when they have high perplexity regarding the context. We also verify that models perform worse in extractive question answering when the perplexity of the answer span in the context is high (see Sec. 3.2). Inspired by these insights, we propose a novel approach named FastMem to improve context awareness via fast memorization of prompts, as depicted in Fig. 1. Specifically, we adapt an instruction fine-tuned model by maximizing the prompt likelihood before generation. To ensure the efficiency of the optimization and prevent overfitting to small text snippets, we update only the last FFN module, achieving significant improvements with minimal computational cost. Our contributions are summarized as follows:
@@ 2 Related Work
Test-Time Training. Test-Time Training (TTT) was initially proposed for vision tasks to adapt models for out-of-distribution data using psuedo-labels (Sun et al., 2020; Bartler et al., 2022). In contrast, FastMem can improve model performance when the train and test data distributions are inconsistent or consistent. Our experiments reveal that a negative correlation between the performance and perplexity with respect to the context exits in LLMs, as discussed in Sec. 3.2. This negative correlation potentially leads to a more general applicability of FastMem. Recently, Hardt and Sun (2024) studied TTT for language models. However, they only consider pretrained models, in which raw text for TTT can be directly fed to the network. If instruction fine-tuned models follow the same method without modification, they risk being corrupted based on our experiments. In FastMem we address this problem with dedicated control tokens. Furthermore, we explore the reduction of latency and achieve real-time application of our method. Thus, FastMem distinguishes itself from existing TTT methods.
@@ 3 Method
In this section, we elaborate on our method, FastMem. In Sec. 3.1, we clarify the notations. In Sec. 3.2, we present the motivation behind our method. In Sec. 3.3, we explain how we develop our method to be robust and efficient. In Sec. 3.4, we explore combining FastMem with existing decoding strategies. Finally, in Sec. 3.5, we demonstrate the use cases. An illustration of our approach and its variants is provided in Fig. 2.
@@ 3.2 Motivation
As introduced in Sec. 1, conflicts between parametric knowledge and contextual information, as well as mass-seeking behavior are supposed to hamper the context awareness. To verify that, we compare the performance of Llama 3-8B-Instruct (Meta AI, 2023) and Qwen 1.5-7B-Chat (Bai et al., 2023) on two datasets: Natural Questions (NQ) (Kwiatkowski et al., 2019) and NQ-SWAP. NQ is constructed based on Wikipedia, which is included in many pre-training corpora. Therefore, the information from NQ has likely been parameterized into Llama 3 and Qwen 1.5. NQ-SWAP is constructed from NQ using the method of Longpre et al. (2021), where we swap the answer strings, which are entities, with random similar entities from the corpus (cf. Sec. B). This process creates new context-answer pairs that the model has not observed before.
We measure the model’s perplexity regarding the minimal span of tokens in the context containing the answer string. As shown in Fig. 3, we observe that models exhibit higher perplexity on NQ-SWAP compared to NQ, proving that the information in NQ belongs to the model’s parametric knowledge. Additionally, the model suffers an obvious utility drop on NQ-SWAP, although the differences between the two datasets are merely short strings involving entities, indicating a negative correlation between perplexity and performance.
The negative correlation suggests that optimizing the model to memorize the context (thereby reducing the perplexity) before inference could be beneficial. This approach can potentially address the commonly hypothesized causes of context-awareness issues, as introduced in Sec. 1. First, reducing the perplexity on context can alleviate potential conflicts between parametric knowledge and contextual information. Second, by increasing the model’s certainty regarding the text in the context, the mass-seeking effect will be reduced.
@@ 3.4 Integrating decoding strategies with FastMem
Li et al. (2023) propose contrastive decoding (CD) which considers the difference between the likelihoods predicted by a strong model and a weak model. Following the concept of CD, we make the prediction of FastMem contrasted with the original prediction made by the model before any memorization took place, as depicted in Fig. 2(b). This is done by subtracting the original prediction probabilities $P_{\bm{\theta}_{0}}(\bm{x}_{t}|\bm{x}_{<t})$ from the current probabilities $P_{\bm{\theta}_{N}}(\bm{x}_{t}|\bm{x}_{<t})$. This subtraction aims to emphasize the differences introduced by the memorization process and mitigate any overfitting by highlighting where the model’s behavior has significantly changed. The modified prediction $\tilde{P}_{\bm{\theta}_{N}}(\bm{x}_{t}|\bm{x}_{<t})$ can be expressed as:
@@
In the second strategy, the final prediction is contrasted with an intermediate representation within the network, following the DoLa framework (Chuang et al., 2024). Specifically, we take the output from an earlier transformer block $k$ where $k<K$, and pass it through the final projection layer and softmax function to obtain an early-exit prediction $P_{\bm{\theta_{0}}^{\{1:k\}}}(\bm{x}_{t}|\bm{x}_{<t})$, as depicted in Fig. 2(c). The final output $P_{\bm{\theta_{N}}}(\bm{x}_{t}|\bm{x}_{<t})$ is then adjusted by subtracting the early-exit prediction. This approach leverages the hierarchical nature of transformer models, where different layers capture different levels of abstraction. By contrasting the final output with an early-exit prediction, we can emphasize the knowledge learned and output by the last fine-tuned FFN. The adjusted prediction can be formalized as:
@@ 4.1 Experimental configurations
Our experiments include evaluations of popular open-source instruction fine-tuned LLMs from the Llama 3 series (8B and 70B) and the Qwen 1.5 series (4B, 7B, and 14B). Additionally, we compare our method with in-context-learning (ICL) and advanced decoding strategies: Decoding by Contrasting Layers (DoLa) (Chuang et al., 2024) and Context-Aware Decoding (CAD) (Shi et al., 2023).
@@
Prompts are carefully crafted and tailored to each dataset and task to elicit optimal performance of the original model, as detailed in Sec. C. For Q&A tasks, we prompt the model to perform extractive answering, while for summarization tasks, we prompt the model to output in required conciseness. Consistent templates are used across models to maintain uniformity in input formatting, with control tokens specialized for different models.
@@ 4.2 Improving context awareness
Tab. 1 presents the performance of various methods using Llama 3-8B-Instruct. Our method consistently outperforms the baseline and is competitive with or superior to decoding strategies across different datasets. In scenarios where there are conflicts between parametric knowledge and contextual information or where the instruction is counter-intuitive for the LLMs’ prior knowledge, our method significantly enhances the baseline model (NQ-SWAP: 59.1% $\rightarrow$ 71.6%, MemoTrap: 77.9% $\rightarrow$ 97.2%). For standard tasks, we find that FastMem is particularly beneficial for complex tasks such as multi-hop reading comprehension (HotpotQA: 67.1% $\rightarrow$ 69.3%). Furthermore, standalone FastMem is robust and generally improves the baseline model across tasks. In contrast, we found that ICL improves performance on some datasets, like Memotrap, but can reduce it on others, such as NQ and HotpotQA. The introduction of ICL examples may include irrelevant information that negatively impacts the model’s performance (although we search for the optimal ICL examples as clarified in Sec. A). CAD performs well on NQ-SWAP and MemoTrap, however, it is detrimental when applied to NQ and HotpotQA.
@@
Tab. 3 demonstrates the effectiveness and robustness of our method across various model architectures and sizes on Q&A tasks. For the Qwen series models, our method consistently enhances performance over the baseline on NQ-SWAP (7B-Chat: 71.9% $\rightarrow$ 75.2%, 14B-Chat: 69.9% $\rightarrow$ 78.4%) and MemoTrap (7B-Chat: 88.8% $\rightarrow$ 98.8%, 14B-Chat: 83.4% $\rightarrow$ 96.7%). We observe that FastMem marginally decreases performance on standard tasks (NQ and HotpotQA) for Qwen 1.5-7B-Chat, but this phenomenon is alleviated for Qwen 1.5-14B-Chat. In contrast, the improvements for Llama 3-70B-Instruct are significant and consistent across all datasets, which have also been observed for Llama 3-8B-Instruct, as shown in Tab. 1. This implies that larger and stronger models can benefit more from our method. In contrast, DoLa has an insignificant impact on zero-shot performance across different models. While CAD performs well on NQ-SWAP and MemoTrap, it often hampers zero-shot performance when applied to NQ and HotpotQA across different models.
@@
In Q&A tasks, we design prompts to guide the model to output structured answers (see Sec. C), thereby preventing the model from blindly increasing accuracy by repeating given contexts. Structured output is also crucial for many applications of LLMs (Liu et al., 2024). However, we find that small-scale models like Qwen 1.5-4B-Chat may produce responses that do not adhere to the template. We mitigate this issue by incorporating format instructions into training, as shown in Fig. 5. Results are provided in Tab. 2. Our approach significantly reduces Qwen 1.5-4B-Chat’s failure rate in producing the required output format (NQ: 33.8% $\rightarrow$ 26.1%, NQ-SWAP: 34.9% $\rightarrow$ 25.5%), thereby greatly improving accuracy. Moreover, we notice that ICL cannot easily address this issue, potentially because ICL examples extend the input length and make it more challenging for small-scale models to understand. Existing contrastive decoding strategies, such as CAD, also cannot address this problem.
@@ 4.3 Ablation study
FastMem is inspired by the negative correlation between perplexity and performance (see Sec. 3.2). To show that performance increases as perplexity decreases, we present the number of epochs vs. performance using Llama 3-8B-Instruct in Fig. 8. As we can see, performance improves as perplexity decreases up to 10 epochs. We also note that after 10 epochs, performance plateaus while perplexity continues to decrease, suggesting that the number of training epochs should not be too many. Additionally, we measured the accuracy and perplexity curves of Llama 3-8B-Instruct and Qwen 1.5-7B-Chat on Memotrap. As shown in Fig. 10 and Fig. 11, the results are consistent with Fig. 8.
@@ 5 Conclusion
6 Acknowledgements
Matthew B. Blaschko and Junyi Zhu received funding from the Flemish Government (AI Research Program) and the Research Foundation - Flanders (FWO) through project number G0G2921N. Tong Xu and Shuochen Liu was supported by the grants from National Natural Science Foundation of China (No.62222213, 62072423).
@@ 7 Limitations
Long-Text Memorization
Llama 3 has a window size of 8K, and the maximum input length is set to 7500 tokens. As the technology of LLMs develops, longer input texts can be expected, which may lead to challenges in fast memorization for real-time applications. Further exploration can be done in this direction.
@@ Appendix A Hyperparameters
We split 20% of the training data to find the optimal ICL configuration. We considered one-shot, two-shot, and three-shot settings, randomly sampling the corresponding number of examples for each and evaluating performance on the holdout dataset. This process was repeated 200 times. We found that the two-shot setting generally performed best across different datasets and models. Therefore, we adopt the two-shot configuration for ICL, using the examples that performed best on the holdout datasets.
@@
Appendix B Construction of NQ-SWAP
NQ-SWAP replaces all mentions of the answer entity in the reference text with other entities to construct new unseen examples. As shown in Fig. 9, we provide an example to help readers understand how NQ-SWAP is constructed based on NQ, following the method outlined by (Longpre et al., 2021).
Given Question: Who has been chosen as the brand ambassador of the campaign ’Beti Bachao, Beti Padhao’? For NQ Reference: On 26 August 2016, Olympic bronze medallist Sakshi Malik was appointed as the brand ambassador for BBBP. Answer: Sakshi Malik For NQ-SWAP Swapped Reference: On 26 August 2016, Olympic bronze medallist Tamara Drasin was appointed as the brand ambassador for BBBP. Swapped Answer: Tamara Drasin
Figure 9: Comparison of NQ and NQ-SWAP.
Appendix C Prompt
Tab. 7 shows the prompts for Llama 3 across all datasets. For models of the Qwen 1.5 series, we simply replace the control tokens. Our prompts are carefully designed to guide the model to respond in alignment with the desired answers. For Llama 3-8B-Instruct, Llama 3-70B-Instruct, Qwen 1.5-7B-Chat, and Qwen 1.5-14B-Chat, we adopt the memorization prompt and optimize the model to memorize the reference, as shown in Fig. 4, in order to improve reading comprehension and summarization performance. To enhance the structured output of Qwen 1.5-4B-Chat, we adopt a memorization prompt that incorporates the instruction segment into memorization, as shown in Fig. 5.